# hdMTD: An R Package for High-Dimensional Mixture Transition Distribution Models

### v1

@@ 5.2 Estimation
With the sample at hand, we proceed to the problem of estimating the relevant lags using the hdMTD functions. The hdMTD package includes four distinct estimators for the set of relevant lags: the BIC, FS, CUT and FSC estimators. An overview of these estimators is provided in Section 3. Users can use either the specified functions hdMTD_"method" (e.g., hdMTD_FS) or the generic hdMTD function and specify the method within its arguments (e.g., method = "FS"). For beginner users, we recommend using the specified functions rather than the generic hdMTD function. This is because the arguments required for the generic hdMTD function depend on the specified method, meaning that the user needs to be familiar with the arguments for the chosen method. In contrast, the specified functions are more user-friendly as their documentation clearly outlines the required arguments for that particular method, making them easier to use.
@@
In the example above, we use the function hdMTD_FS to obtain, within the first 40 pasts (hence d = 40), those l = 4 pasts that are most relevant according to the FS estimator [^4] The FS estimator initially identifies the past that has the most significant predictive power for the present. Then, if $l>1$, it iteratively searches for the lag that is most important given the knowledge of the previous lags, and so on. In our output, the lag -1 was found to be the most relevant, followed by lag -30 in conjunction with the information from lag -1, and so forth.. Observe that the same result is obtained by using hdMTD(X, d = 40, method = "FS", l = 4). The function hdMTD_FS was capable of retrieving the relevant lags 30-30, 15-15 and 1-1. Since l = 4, it also brought an additional lag, the 27-27 in the example. Had we used l = 3, the function would have returned exactly Lambda, given that $27$ was the last output.
Now, let us see how the other estimators work. We can look for the 4 relevant pasts of the MTD model that minimizes (8) with the hdMTD_BIC function.
@@
[1]  1 15 17 30 In this example, the BIC method computes the penalized log-likelihood of all possible models with 4 relevant pasts between 1 and 40 (there are $\binom{40}{4}$ such models). Among all combinations of 4 lags, the combination of lags 30,17,15,1-30,-17,-15,-1 has the smallest penalized log-likelihood. Note that from this output alone, we cannot determine which of the 4 returned lags is more important (differently from the output of hdMTD_FS).
@@
[1]  1 15 17 30 Now hdMTD_BIC searched only across all subsets of S of size 4 (since minl=maxl=4). Given that $\{1,15,17,30\}\subset S$ has size 4, the output is the same as in the unconstrained case. However, the computation time is much faster (less than 10 seconds). We can also allow the function to look for sets with other sizes by changing minl and maxl arguments.
@@
[1]  30 The hdMTD_BIC function iterates over all possible subsets with size 1 (in this case, there were 10 of them), then over all possible subsets of size 2, and so on, until reaching all possible subsets of size 4 (since minl = 1, maxl = 4). In this scenario, the model with single relevant past 30-30 is the most likely given the sample and the penalization criterion (recall that BIC penalizes the number of parameters). Since the model with pasts 30-30, 17-17, 15-15, and 1-1 was within the possibilities, this shows that, given the sample, the BIC criterion selects a more parsimonious model, for the increase in likelihood did not outweigh the loss from the extra parameters. The penalization term is multiplied by a constant which can be passed to the model with the argument xi, by default, xi = 0.5. Changing this constant may alter the output, as a larger constant increases the penalty term, while a smaller constant decreases it. The hdMTD_BIC function can return, for each size (i.e., from size minl to maxl), the most likely set, by setting the argument byl = TRUE. Setting BICvalue to TRUE, makes the function return the BIC value of each selected set.
@@
      30    15,30   1,15,30  1,15,17,30 smallest: 1,15,17,30641.7328 643.1757  642.5873  641.3069               641.3069 With this change in the constant, the penalty across more parameters was reduced and the model with four lags returned the smallest BIC value.
@@
[1] 40 35 30 27 20 17 15 10  5  1 In this case, no past was cut, which means our threshold is too low. We can increase it by raising alpha, for example:
@@
[1] 35 27  5  1 Just like our first use of the hdMTD_BIC function, we could have left the argument S empty and the function would set S=1:d. However, the CUT method verifies the relevance of each lag while taking into account all other lags in S. Not reducing our search to a subset S would make the function hdMTD_CUT work with sequences of size d (in this case, $40$) at all times, increasing significantly the running time. For example, as it is in the above example (with 10 elements in S), the function hdMTD_CUT took about 3 minutes to run. But if we add two more lags to S (for example S = c(1,2,3,5,10,15,17,20,27,30,35,40)), the process can take about 35 minutes. Besides, the elements in S change the probabilities estimated within the function and might alter its output. In the example below, even though lags 35-35 and 27-27 were potential choices, when we reduce S, the output changes.
@@
[1] 30 17  5  1 Finally, the FSC estimator combines the FS and CUT estimators. It uses the function hdMTD_FS to reduce the set 1:d to some subset S of size l, and then calls hdMTD_CUT to trim out the less important pasts. However, there is a distinction between employing the FS estimator followed by the CUT estimator and using the FSC estimator. In the FSC estimator, the sample is divided in half. The first half is used to compute the FS estimator, while the CUT estimator is computed from the second half. This is done to ensure that the CUT estimator is computed accordingly to the theory developed in Ost and Takahashi (2023).
@@
[1] 30 24 If you compare the output of the FS estimator with this output from FSC estimator, you might wonder about the appearance of the past 24-24 (since hdMTD_FS(X, d = 40, l = 4) returned 30, 15, 1, 27). You need to take into account the splitting of the sample. In this example, the input set the FS estimator provided to FSC estimator was:
@@
[1] 11 30  7 24 Afterwards, the CUT estimator uses the second half of the sample to remove pasts 11-11 and 7-7. To conclude, all arguments in the FSC estimator are exactly those used in FS and CUT estimators, and the order of the output no longer reflects the importance of each lag, as it did in FS. In fact, FS is the only estimator where the order of the output carries informative value.
Now let us assume we have estimated a set of relevant pasts $\hat{\Lambda}$ from the data. If we wish to estimate the transition matrix P from the sample we can set S $\hat{\Lambda}$ and call the probs function.
R> probs(X, S = c(1, 15, 30))
@@
1                000   0 0.5000000
2                000   1 0.5000000
3                001   0 0.3714286
4                001   1 0.6285714
5                010   0 0.6306306
6                010   1 0.3693694
7                011   0 0.5065789
8                011   1 0.4934211
9                100   0 0.3861386
10               100   1 0.6138614
11               101   0 0.1987952
12               101   1 0.8012048
13               110   0 0.3888889
14               110   1 0.6111111
15               111   0 0.3697917
16               111   1 0.6302083
R> probs(X,S=c(1,15,30),matrixform=TRUE)
            0         1
000 0.5000000 0.5000000
001 0.3714286 0.6285714
010 0.6306306 0.3693694
011 0.5065789 0.4934211
100 0.3861386 0.6138614
101 0.1987952 0.8012048
110 0.3888889 0.6111111
111 0.3697917 0.6302083
When using the argument matrixform = TRUE, the output is formatted into a matrix. In matrix form, for example, the row $110$ displays the estimated transition probabilities when $x_{-30}=1$, $x_{-15}=1$ and $x_{-1}=0$ (as can be seen in the data frame output).
@@
     -1        -15        -300.1233648 0.1017517 0.1846987 This function can also estimate the oscillations if you provide it with a sample of a MTD model and a set of lags $S$ (that must be labeled as S). In this case the function will assume that d = max(S) and estimate the quantities in (3) for each lag in S.
@@
The package also offers a method for estimating all MTD parameters using the expectation maximization algorithm (EM) via the MTDest function. This function was developed based on the algorithm proposed by Lèbre and Bourguignon (2008). It iteratively updates parameters to maximize likelihood until no significant increase is observed or a predetermined number of iterations is reached. To execute this procedure, an initial list of parameters must be provided through the init argument.
@@
R> init <- list(+       ’lambdas’ = c(0.01, 0.33, 0.33, 0.33),+       ’p0’ = c(0.5, 0.5),+       ’pj’ = rep(list(matrix(c(0.5, 0.5, 0.5, 0.5), ncol = 2, nrow = 2)), 3)+       ) This list must have the following entries: a vector named lambdas as the MTD weights; a list of matrices named pj as the MTD past conditional distributions. If your MTD has an independent distribution the list should also have a vector called p0 with the independent distribution. Moreover, if your MTD does not have a $p_{0}$ distribution you must set the first entry in lambdas to 0. The MTDest parameters update will halt by default if either the likelihood increase at some iteration becomes lower than M = 0.01 or if the number of iterations nIter reaches 100. Both these arguments can be modified by the user.
R> MTDest(X, S = c(1, 15, 30), init = init, iter = TRUE)
\$lambdas      lam-0       lam-1      lam-15      lam-300.009353147 0.323151043 0.326490125 0.341005685\$pj\$pj\$‘p_-1‘          0         10 0.5982386 0.40176141 0.2641061 0.7358939\$pj\$‘p_-15‘          0         10 0.1909446 0.80905541 0.5495934 0.4504066\$pj\$‘p_-30‘          0         10 0.6978307 0.30216931 0.2113007 0.7886993\$p0   p_0(0)    p_0(1)0.3911887 0.6088113\$iterations[1] 9\$distlogL [1] 36.20577891  7.20855924  2.95676388  1.22724939  0.51347791  0.21602936 [7]  0.09125854  0.03867624  0.01643732  0.00700394
@@
In the output, lambdas, pj, and p0 are the updated estimated parameters of our MTD model. Iterations and distlogL are included because we have set the argument iter to TRUE in the function (by default iter = FALSE). Iterations denotes the number of times the model parameters (provided through init) were updated. distlogL computes the distance of log-likelihoods between updated and non-updated parameters at each iteration. As soon as this distance becomes lower than $0.01$ (because M = 0.01 by default), the updates stops. It is also possible to set M to NULL, and the function MTDest will only halt when the maximum number of iterations nIter is reached (by default nIter = 100). Additionally, there is the argument oscillations that can be set to TRUE if the user wishes the function to calculate the oscillations based on the outputted parameters.
@@
R> MTDest(X, S=c(1, 15, 30), M = NULL, nIter = 9, init = init,+       oscillations = TRUE)
@@
\$lambdas      lam-0       lam-1      lam-15      lam-300.009353147 0.323151043 0.326490125 0.341005685\$pj\$pj\$‘p_-1‘          0         10 0.5982386 0.40176141 0.2641061 0.7358939\$pj\$‘p_-15‘          0         10 0.1909446 0.80905541 0.5495934 0.4504066\$pj\$‘p_-30‘          0         10 0.6978307 0.30216931 0.2113007 0.7886993\$p0   p_0(0)    p_0(1)0.3911887 0.6088113\$oscillations       -1       -15       -300.1065876 0.1159706 0.1643166 In this example, the function MTDest stopped after $9$ updates because we have set nIter = 9. The estimated parameters in the output should be exactly the same as before, as there were the same number of iterations.
If the user wishes to estimate the transition matrix based on the outputs of the MTDest function, resort to the already mentioned function MTDmodel.
R> estParam <- MTDest(X, S = c(1, 15, 30), init = init)R> estMTD <- MTDmodel(Lambda, A, lam0 = estParam\$lambdas[1],+       lamj = estParam\$lambdas[-1], p0 = estParam\$p0, pj = estParam\$pj)R> estMTD\$P
@@ 5.3 Testing hdMTD
R> set.seed(123)R> Lambda <- c(1, 5)R> A <- c(0, 1)R> lam0 <- 0.01R> p0 <- c(0.5, 0.5)R> MTD <- MTDmodel(Lambda, A, lam0, p0 = p0) we generate $N_{rep}=100$ replications of size N = 10000 using the perfectSample function.
@@
$\bar{\Delta}_{\mathrm{FS}}(m)$ 0.02847 0.02026 0.01752 0.01561 0.01442 0.01010 0.00681
$\bar{\Delta}_{\mathrm{Oracle}}(m)$ 0.00728 0.00916 0.01104 0.01167 0.01150 0.00971 0.00681
$\bar{\Delta}_{\mathrm{Naive},5}(m)$ 0.06786 0.05845 0.05534 0.04974 0.04485 0.03412 0.02508
$\bar{\Delta}^{\mathrm{std}}_{\mathrm{FS}}(m)$ 0.06469 0.04605 0.03981 0.03547 0.03278 0.02295 0.01547
$\bar{\Delta}^{\mathrm{std}}_{\mathrm{Oracle}}(m)$ 0.01655 0.02081 0.02509 0.02651 0.02614 0.02207 0.01547
$\bar{\Delta}^{\mathrm{std}}_{\mathrm{Naive},5}(m)$ 0.15420 0.13281 0.12577 0.11303 0.10193 0.07753 0.05700
@@
As can be seen, the mean error of the FS estimator seems to converge to that of the Oracle of size 2 estimator as $m$ increases. In fact, when using the full realizations (i.e., $m=10000$), the FS method and the Oracle of size 2 agreed that the relevant lag set was $\{-5,-1\}$ across all $100$ replications. For $m=5000$, they only disagreed in 5 replications. This agreement reflects the high precision and consistency of the FS estimator as $m$ grows. In contrast, the Naive estimator —which has access to the true relevant lags — exhibited substantially higher errors in total variation distance to $\operatorname{\mathsf{P}}(\cdot|00_{\Lambda})$ across all values of $m$. This suggests that simply including the correct lags is not sufficient: without appropriate model selection the estimator may fail to capture the true conditional distribution effectively, especially for smaller sample sizes.
@@
The hdMTD package includes the tempdata dataset, a tibble containing hourly temperature measurements from Brasília, Brazil. The dataset comprises three columns: DATE (ranging from January 01, 2003 to August 31, 2024), TIME (hourly intervals from 00:00 to 23:00 in UTC-3), and MAXTEMP (maximum temperature in Celsius recorded within each hour interval). With 189,936 observations spanning 7,914 complete days, this dataset provides comprehensive coverage of Brasília’s thermal patterns at 15.78°S, 47.92°W (altitude: 1,159.54m). The data was collected by Brazil’s National Institute of Meteorology (National Institute of Meteorology, 2024) and can be loaded in R using data("tempdata").
@@
# A tibble: 4 × 3  DATE       MAXTEMP MAXTEMP1  <date>       <dbl>    <dbl>1 2010-08-05    20.7        12 2010-08-06    20.5        13 2010-08-07    21.8        24 2010-08-08    22.3        2 The categorization scheme is as follows:
@@
R> ndays <- nrow(temp %>% filter(DATE >= "2023-09-01"))R> Temp12_Train <- Temp12[-seq_len(ndays)]R> Temp12_Test <- Temp12[seq_len(ndays)] We rerun the hdMTD_FS function on Temp12_Train to account for the reduced sample size. Interestingly, the resulting set of selected lags remains unchanged (1-1, 364-364, and 6-6), indicating that these dependencies are stable and persist even when the last year of data is excluded.

### v2

@@ 5.2 Estimation
With the sample at hand, we proceed to the problem of estimating the relevant lags using the hdMTD functions. The hdMTD package includes four distinct estimators for the set of relevant lags: the BIC, FS, CUT and FSC estimators. An overview of these estimators is provided in Section 3. Users can either use the specific functions hdMTD_"method" (e.g., hdMTD_FSC()) or the more general hdMTD() function, indicating the method via the method argument (e.g., method = "FSC"). The specific functions provide a more direct interface, with tailored documentation and argument prompts that guide the user more effectively. The hdMTD() function, on the other hand, returns a structured object with additional information, which can be useful for further inspection. However, note that the arguments required by hdMTD() depend on the chosen method and are not directly visible.
@@
In the example above, we use the function hdMTD_FS to obtain, within the first 40 pasts (hence d = 40), those l = 4 pasts that are most relevant according to the FS estimator [^4] The FS estimator initially identifies the past that has the most significant predictive power for the present. Then, if $l>1$, it iteratively searches for the lag that is most important given the knowledge of the previous lags, and so on. In our output, the lag -1 was found to be the most relevant, followed by lag -30 in conjunction with the information from lag -1, and so forth..
Alternatively, the same estimation can be performed using the general hdMTD() function with method = "FS", which returns an object with additional information:
R> FS <- hdMTD(X, d = 40, method = "FS", l = 4)R> S(FS); summary(FS)
[1] 30 15 1 27hdMTD lag selectionCall:hdMTD(X = X, d = 40, method = "FS", l = 4)Method: FSOrder upper bound (d): 40Selected S set: 30, 15, 1, 27Relevant lag set estimated by FS method : -30, -15, -1, -27
The hdMTD() function returns an object of class hdMTD, which stores the estimated lags along with auxiliary information. The selected lags can be accessed using the accessor S() (output in $\mathbb{Z}^{+}$) or lags() (output in $\mathbb{Z}^{-}$), and additional details are available via the summary(). To include the full list of arguments used during estimation, use summary(x, settings = TRUE). In the example above, the estimated set contains the true lags 30-30, 15-15, and 1-1. Since l = 4, the algorithm also included one additional lag (27-27 in this case). Had we used l = 3, the function would have returned exactly Lambda, given that $27$ was the last output.
Now, let us see how the other estimators work. In what follows, we illustrate each method using its specific function (e.g., hdMTD_BIC), which provides a direct output of the selected lags. Alternatively, all estimators can also be accessed through the general hdMTD() interface by specifying the corresponding method. We can look for the 4 relevant pasts of the MTD model that minimizes (8) with the hdMTD_BIC function.
@@
[1] 1 15 17 30
In this example, the BIC method computes the penalized log-likelihood of all possible models with 4 relevant pasts between 1 and 40 (there are $\binom{40}{4}$ such models). Among all combinations of 4 lags, the combination of lags 30,17,15,1-30,-17,-15,-1 has the smallest penalized log-likelihood. Note that from this output alone, we cannot determine which of the 4 returned lags is more important (differently from the output of hdMTD_FS).
@@
[1] 1 15 17 30
Now hdMTD_BIC searched only across all subsets of S of size 4 (since minl=maxl=4). Given that $\{1,15,17,30\}\subset S$ has size 4, the output is the same as in the unconstrained case. However, the computation time is much faster (less than 10 seconds). We can also allow the function to look for sets with other sizes by changing minl and maxl arguments.
@@
[1] 30
The hdMTD_BIC function iterates over all possible subsets with size 1 (in this case, there were 10 of them), then over all possible subsets of size 2, and so on, until reaching all possible subsets of size 4 (since minl = 1, maxl = 4). In this scenario, the model with single relevant past 30-30 is the most likely given the sample and the penalization criterion (recall that BIC penalizes the number of parameters). Since the model with pasts 30-30, 17-17, 15-15, and 1-1 was within the possibilities, this shows that, given the sample, the BIC criterion selects a more parsimonious model, for the increase in likelihood did not outweigh the loss from the extra parameters. The penalization term is multiplied by a constant which can be passed to the model with the argument xi, by default, xi = 0.5. Changing this constant may alter the output, as a larger constant increases the penalty term, while a smaller constant decreases it. The hdMTD_BIC function can return, for each size (i.e., from size minl to maxl), the most likely set, by setting the argument byl = TRUE. Setting BICvalue to TRUE, makes the function return the BIC value of each selected set.
@@
30 15,30 1,15,30 1,15,17,30 smallest: 1,15,17,30641.7328 643.1757 642.5873 641.3069 641.3069
With this change in the constant, the penalty across more parameters was reduced and the model with four lags returned the smallest BIC value.
@@
[1] 1 5 10 15 17 20 27 30 35 40
In this case, no past was cut, which means our threshold is too low. We can increase it by raising alpha, for example:
@@
[1] 1 5 27 35
Just like our first use of the hdMTD_BIC function, we could have left the argument S empty and the function would set S=1:d. However, the CUT method verifies the relevance of each lag while taking into account all other lags in S. Not reducing our search to a subset S would make the function hdMTD_CUT work with sequences of size d (in this case, $40$) at all times, increasing significantly the running time. For example, as it is in the above example (with 10 elements in S), the function hdMTD_CUT took about 3 minutes to run. But if we add two more lags to S (for example S = c(1,2,3,5,10,15,17,20,27,30,35,40)), the process can take about 35 minutes. Besides, the elements in S change the probabilities estimated within the function and might alter its output. In the example below, even though lags 35-35 and 27-27 were potential choices, when we reduce S, the output changes.
@@
[1] 1 5 17 30
Finally, the FSC estimator combines the FS and CUT estimators. It uses the function hdMTD_FS to reduce the set 1:d to some subset S of size l, and then calls hdMTD_CUT to trim out the less important pasts. However, there is a distinction between employing the FS estimator followed by the CUT estimator and using the FSC estimator. In the FSC estimator, the sample is divided in half. The first half is used to compute the FS estimator, while the CUT estimator is computed from the second half. This is done to ensure that the CUT estimator is computed accordingly to the theory developed in Ost&Takahashi.
@@
[1] 24 30
If you compare the output of the FS estimator with this output from FSC estimator, you might wonder about the appearance of the past 24-24 (since hdMTD_FS(X, d = 40, l = 4) returned 30, 15, 1, 27). You need to take into account the splitting of the sample. In this example, the input set the FS estimator provided to FSC estimator was:
@@
[1] 11 30 7 24
Afterwards, the CUT estimator uses the second half of the sample to remove lags 11-11 and 7-7. To conclude, all arguments in the FSC estimator are exactly those used in FS and CUT estimators, and the order of the output no longer reflects the importance of each lag, as it did in FS. In fact, FS is the only estimator where the order of the output carries informative value.
Now let us assume we have estimated a set of relevant lags $\hat{\Lambda}$ from the data. If we wish to estimate the transition matrix P from the sample we can set S $\hat{\Lambda}$ and call the empirical_probs function.
R> head(empirical_probs(X, S = c(1, 15, 30)), 6)
@@
R> empirical_probs(X, S = c(1, 15, 30), matrixform = TRUE)
0 1000 0.5000000 0.5000000001 0.3714286 0.6285714010 0.6306306 0.3693694011 0.5065789 0.4934211100 0.3861386 0.6138614101 0.1987952 0.8012048110 0.3888889 0.6111111111 0.3697917 0.6302083
When using the argument matrixform = TRUE, the output is displayed as a matrix. As a reading guide for transition matrices in the hdMTD package, rows represent past contexts (read from oldest to most recent). For example, the row 001, in this case, gives the estimated transition probabilities given $x_{-30}=0$, $x_{-15}=0$, and $x_{-1}=1$ (as can be seen by comparing with the data frame output).
@@
-1 -15 -300.1233648 0.1017517 0.1846987
This function can also estimate the oscillations if you provide it with a sample of a MTD model and a set of lags $S$. In this case, the function will assume that d = max(S) and estimate the quantities in (3) for each lag in S.
@@
The package also offers a method for estimating all MTD parameters using the expectation maximization algorithm (EM) via the MTDest function. This function was developed based on the algorithm proposed by EMforMTD and produces an S3 object of class MTDest. It iteratively updates parameters to maximize log-likelihood until convergence or until a predetermined number of iterations is reached. To execute this procedure, an initial list of parameters must be provided through the init argument:
R> init <- list(+ ’lambdas’ = c(0.01, 0.33, 0.33, 0.33),+ ’p0’ = c(0.5, 0.5),+ ’pj’ = rep(list(matrix(c(0.5, 0.5, 0.5, 0.5), ncol = 2, nrow = 2)), 3)+ )
@@
This list must have the following entries: a vector named lambdas with the MTD weights; a list of matrices named pj containing the MTD past conditional distributions. If your MTD has an independent distribution the list should also have a vector called p0 with the independent distribution. If no independent distribution is present, the first entry of lambdas must be set to zero. The MTDest parameters update will halt by default if either the likelihood increase at some iteration becomes smaller than M = 0.01 or if the number of iterations nIter reaches 100. Both these arguments can be modified by the user.
R> emMTD <- MTDest(X, S = c(1, 15, 30), init = init, iter = TRUE)R> summary(emMTD)
Summary of EM estimation for MTD model:Call:MTDest(X = X, S = c(1, 15, 30), init = init, iter = TRUE)Lags (-S): -1, -15, -30State space (A): 0, 1lambdas (weights): lam0 lam-1 lam-15 lam-300.009353147 0.323151043 0.326490125 0.341005685Independent distribution p0: p0(0) p0(1)0.3911887 0.6088113Transition matrices pj (one per lag): pj for lag j = -1: 0 10 0.5982386 0.40176141 0.2641061 0.7358939 pj for lag j = -15: 0 10 0.1909446 0.80905541 0.5495934 0.4504066 pj for lag j = -30: 0 10 0.6978307 0.30216931 0.2113007 0.7886993Log-likelihood: -623.879Iterations Report:Number of updates: 9Last compared difference of logLik: 0.00700394
In the output, lambdas, pj, and p0 are the updated estimated parameters of the MTD model. Since iter = TRUE, the summary() also reports the iteration diagnostics, confirming convergence when the last log-likelihood difference dropped below M = 0.01. The fitted object additionally stores the log-likelihood of the model, which is displayed in the summary() output and can be directly accessed with the logLik() method. All available methods and accessors for MTDest objects are listed in the print() method output.
@@
It is also possible to set M to NULL, and the function MTDest will only halt when the maximum number of iterations nIter is reached (by default nIter = 100). Additionally, there is the argument oscillations that can be set to TRUE if the user wishes the function to calculate the oscillations based on the outputted parameters.
@@
R> emMTD <- MTDest(X, S = c(1, 15, 30), M = NULL, nIter = 9,+ init = init, oscillations = TRUE)R> summary(emMTD)
Summary of EM estimation for MTD model:Call:MTDest(X = X, S = c(1, 15, 30), M = NULL, init = init, nIter = 9, oscillations = TRUE)Lags (-S): -1, -15, -30State space (A): 0, 1lambdas (weights): lam0 lam-1 lam-15 lam-300.009353147 0.323151043 0.326490125 0.341005685Independent distribution p0: p0(0) p0(1)0.3911887 0.6088113Transition matrices pj (one per lag): pj for lag j = -1: 0 10 0.5982386 0.40176141 0.2641061 0.7358939 pj for lag j = -15: 0 10 0.1909446 0.80905541 0.5495934 0.4504066 pj for lag j = -30: 0 10 0.6978307 0.30216931 0.2113007 0.7886993Log-likelihood: -623.879Oscillations: -1 -15 -300.1079753 0.1170953 0.1659095
@@
In this example, the function MTDest stopped after $9$ updates because we have set nIter = 9. The estimated parameters in the output should be exactly the same as before, as there were the same number of iterations.
Since MTDest objects do not carry a global transition matrix the transitP() accessor is not available for it. However, the user can coerce an MTDest object to an MTD via the as.MTD() function and thus access any methods for it:
R> emMTD <- MTDest(X, S = c(1, 15, 30), init = init)R> MTD_hat <- as.MTD(emMTD)R> transitP(MTD_hat)
@@ 5.3 Testing hdMTD
R> set.seed(123)R> Lambda <- c(1, 5)R> A <- c(0, 1)R> lam0 <- 0.01R> p0 <- c(0.5, 0.5)R> MTD <- MTDmodel(Lambda, A, lam0, p0 = p0)
we generate $N_{rep}=100$ replications of size N = 10000 using the perfectSample function.
@@
$\bar{\Delta}_{\mathrm{FS}}(m)$ 0.03279 0.01940 0.01686 0.01492 0.01438 0.00980 0.00649
$\bar{\Delta}_{\mathrm{Oracle}}(m)$ 0.00844 0.01006 0.01201 0.01143 0.01194 0.00971 0.00649
$\bar{\Delta}_{\mathrm{Naive},5}(m)$ 0.06898 0.05988 0.05072 0.04862 0.04381 0.03072 0.02330
$\bar{\Delta}^{\mathrm{std}}_{\mathrm{FS}}(m)$ 0.07452 0.04410 0.03832 0.03390 0.03268 0.02228 0.01475
$\bar{\Delta}^{\mathrm{std}}_{\mathrm{Oracle}}(m)$ 0.01917 0.02287 0.02729 0.02598 0.02714 0.02207 0.01475
$\bar{\Delta}^{\mathrm{std}}_{\mathrm{Naive},5}(m)$ 0.15676 0.13608 0.11525 0.11049 0.09957 0.06980 0.05296
@@
As can be seen, the mean error of the FS estimator seems to converge to that of the Oracle of size 2 estimator as $m$ increases. In fact, when using the full realizations (i.e., $m=10000$), the FS method and the Oracle of size 2 agreed that the relevant lag set was $\{-5,-1\}$ across all $100$ replications. For $m=5000$, they only disagreed in 3 replications. This agreement reflects the high precision and consistency of the FS estimator as $m$ grows. In contrast, the Naive estimator —which has access to the true relevant lags — exhibited substantially higher errors in total variation distance to $\operatorname{\mathsf{P}}(\cdot|00_{\Lambda})$ across all values of $m$. This suggests that simply including the correct lags is not sufficient: without appropriate model selection the estimator may fail to capture the true conditional distribution effectively, especially for smaller sample sizes.
@@
The hdMTD package includes the tempdata dataset, a tibble containing hourly temperature measurements from Brasília, Brazil. The dataset comprises three columns: DATE (ranging from January 01, 2003 to August 31, 2024), TIME (hourly intervals from 00:00 to 23:00 in UTC-3), and MAXTEMP (maximum temperature in Celsius recorded within each hour interval). With 189,936 observations spanning 7,914 complete days, this dataset provides comprehensive coverage of Brasília’s thermal patterns at 15.78°S, 47.92°W (altitude: 1,159.54m). The data was collected by Brazil’s National Institute of Meteorology (inmet2024) and can be loaded in R using data("tempdata").
@@
# A tibble: 4 × 3 DATE MAXTEMP MAXTEMP1 <date> <dbl> <dbl>1 2010-08-05 20.7 12 2010-08-06 20.5 13 2010-08-07 21.8 24 2010-08-08 22.3 2
The categorization scheme is as follows:
@@
R> ndays <- nrow(temp %>% filter(DATE >= "2023-09-01"))R> Temp12_Train <- Temp12[-seq_len(ndays)]R> Temp12_Test <- Temp12[seq_len(ndays)]
We rerun the hdMTD_FS function on Temp12_Train to account for the reduced sample size. Interestingly, the resulting set of selected lags remains unchanged (1-1, 364-364, and 6-6), indicating that these dependencies are stable and persist even when the last year of data is excluded.